OpenStudy (anonymous):

does cos^(2)n/(2^n) converge?

OpenStudy (zarkon):

are you asking if \(\displaystyle \lim_{n\to\infty}\frac{\cos^2(n)}{2^n}\) converges?

OpenStudy (anonymous):

It does

OpenStudy (anonymous):

how did you come to that conclusion? Can you show me the math?

OpenStudy (anonymous):

\[-1 \leq\cos^2(n) \leq 1\] \[\frac{-1}{2^n} \leq\frac{\cos^2(n)}{2^n}\leq\frac{1}{2^n}\] evaluating both the limits as n goes to infinity of -1/2^n and 1/2^n gives us 0 for both\[0\leq\lim_{n\rightarrow \infty} \frac{\cos^2(n)}{2^n} \leq 0\]applying the squeeze theorem here will show you that the sequence converges.

OpenStudy (anonymous):

thanks!